# Implemetation of a shooting technique for quantum optimal control on spin qudits

### 23 Mar 2026 v1

@@ IV Conclusion
In this article, we introduced an optimization method to shape pulses for implementing quantum gates on spin qudits systems. We demonstrated its effectiveness relatively to implementation times for medium to large spin systems, the interest of the method growing with the Hilbert space size. The method was showcased on two SMM, the $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$ molecules, taking into account some of their specific constraints.
@@
    Experimentally implementing the MAGICARP on a system to see its advantages in action.
@@
    Having a better understanding of the optimization landscape of the MAGICARP to improve its convergence speed.
@@
    Understand which gradient descent methods could power up the MAGICARP.
@@
    Getting precise insights on how the connectivity between levels of the Hilbert space the MAGICARP is acting on can influence its efficiency.
  5. 5.
    Seeing how the MAGICARP can approach the quantum speed limit.
@@ A.1 Importance of the graph
The graph on which simulations are taken are ones like the double decker from fig. 2a, linearly coupled graphs, and the one of the triple decker from fig. 2b. If we assume different graphs, possible in other physical molecules and platforms, we could have different time implementations for the same target gates and Hilbert space dimension. A most simple configuration is the linear graph where each level is only coupled to its nearest neighbours. It’s the congifuration of the fig. 1a. Another extremal configuration is the complete graph where each level is coupled to all the others. We can test how the GRAPE method and the MAGICARP take advantage of the growing connectivity of the graph. We will plot how the MAGICARP and the GRAPE method compare to themselves when runned over a complete graph rather than on a linearly coupled one in fig. 9. Note how in both case of fig. 9b the smallest implementation time found is increasingly improved by the fact of using a complete graph instead of a linear one (thus having more available controls) with the size of the Hilbert space. In the case of fig. 9a the MAGICARP is slightly helped by the use of a complete graph instead of a linear one (gates are in median around 20% faster in all dimension) but is not as much of an improvement as for GRAPE. Thus sparsely connected graph seem already convenient for the MAGICARP.
Figure 9: Figure comparing the median or the minimum time to make at least 100 gates with a precision of 1e-4 for the MAGICARP and the GRAPE method on complete graphs with 3 to 9 nodes with themselves on linear graphs. For each optimization method, the quantity shown is the ratio between the runtime on a linear graph and the runtime on a complete graph, effectively describing how each method is faster when being run on a complete graph instead of a linear one. (a) Ratio of the median execution time for 100 random Haar gates. (b) Ratio of the minimum execution time for a QFT gate.
@@ A.2 Importance of a $\sigma_{z}$ control
Figure 10: Figure comparing the median or the minimum time to make at least 100 gates with a precision of 1e-4 for the MAGICARP and the GRAPE method on complete and linear graphs with 3 to 9 nodes with themselves when allowing for a control on $\sigma_{z}$. For each optimization method, the quantity shown is the ratio between the runtime with only $\sigma_{x}$ and $\sigma_{y}$ controls and the runtime with all three controls on each edge of the graph the method has been run on. It effectively describes how each situation (method + graph it is run on) is faster when being run while allowing a control on $\sigma_{z}$ instead of just $\sigma_{x}$ and $\sigma_{y}$. The red curves correspond to the aforementioned ratio for the GRAPE method run on a linear graph, the orange ones for the GRAPE method run on a complete graph, the green ones for the MAGICARP run on a lienar graph, and the cyan ones for the MAGICARP run on a complete graph. The x-axis describes the Hilbert space dimension the optimization methods have been run on. (a) Ratio of the median execution time for 100 random Haar gates. (b) Ratio of the minimum execution time for a QFT gate.
One can note that adding a control on $\sigma_{z}$ is not improving much the median solutions found on random gates in fig. 10a or the minimum solutions found in fig. 10b for the specific QFT gate in all situations but in the case of the MAGICARP method run on a complete graph, the cyan curve of fig. 10a, especially when the dimension grows. It applies only for the median implementation time since the same cyan curve in fig. 10a is not affected. It would be interesting to understand further why adding a $\sigma_{z}$ control gives a particular advantage to the MAGICARP on a strongly connected graph.
@@ B.1 Improvement of the GRAPE with the target fidelity
The GRAPE method gives better implementation times according to the target fidelity it has too reach.
@@
It is understandable since being able to be further away from the target obviously relieves a constraint on the pulses. We chose to compare implementation times at 1e-4 of target fidelity for the GRAPE when comparing to other algorithms for multiple reasons. First, this is already a fidelity allowing for making a short algorithm without being bothered by this lack of precision of the gates while in many platforms noise is already inducing an error of this order [25]. Second, Maintaining the fidelities of pulses found by the MAGICARP to be strictly lower than 1e-4 proved difficult with the gradient we used. Third, at lower fidelities than 1e-4 GRAPE is not giving significantly slower pulses than for this value of the target fidelity as shown infig. 11.
@@ B.2 Improvement of the GRAPE with the number of controls
The GRAPE method gives better implementation times according to the number of piece-wise amplitude steps it can use, though up to a limit. In all this article, we chose to compare implementation times with $10(\text{dim}+1)$ piece-wise amplitude steps for each controls, with dim being the dimension of the Hilbert space at hand. It is a large enough number of amplitude to not see improvements by taking a larger amount of piece-wise amplitude steps for each controls as shown infig. 12.
@@ Appendix C Numerical framework
Our objectives in this section is twofold. We provide closed-form expressions for the derivative of the map that associates to every control parameter $g$ the final propagator of Schrödinger’s equation. We outline its derivation which relies on the so-called adjoint state method. This provide the tools needed to implement any first-order optimization routine and possibly to adapt the $g$ parametrization of controls to another problem. On the other hand, we propose our implementation of the natural gradient descent, an optimization routine used to produce some of the figures in this paper. Julia and Python codes are available on the Github repositories Magicarp and pyMagicarp.
@@ C.1 Two equivalent optimal control problems in the driftless case
The third Assumption C.3 is common in the practical coherent control of quantum systems because the control Hamiltonians often do not linearly span the set of tangent vectors to the special unitary group at the identity matrix. The Lie algebra rank condition is about iterated commutators of the control Hamiltonians. When it holds combining the available control directions provides a family of tangent vectors that is sufficiently rich to move in any admissible direction. Intuitively this is because to second order in $t$ close to zero, $e^{-tA}e^{-tB}e^{tA}e^{tB}=I+t^{2}[A,B]+o(t^{2})$. All in all Assumption C.3 is meant to guarantee that any desired special unitary gate is reachable by means of an appropriate sequence of constant pulses.
@@
Let $P$ stand for the orthogonal projector onto the linear span $\mathscr{H}_{0}\subset\operatorname{i}\mathfrak{su}(d)$ of $H_{1},\ldots,H_{m}$ for the norm $g\mapsto\operatorname{Tr}(gg^{\dagger})^{1/2}$. Consider the following two optimal control problems. On the one hand
@@
subject to the pointwise constraint $\sum_{j=1}^{m}u_{j}(t)^{2}\leq 1$ and to the terminal constraint $U(T)=U_{f}$, where the propagator $U$ solves the Cauchy problem $U(0)=I$ and $\dot{U}=-\operatorname{i}\sum_{j=1}^{m}u_{j}(t)H_{j}U$ for $0<t<T$.
@@
$\inf_{g\in\operatorname{i}\mathfrak{su}(d)}|Pg|$ ($\mathscr{P}_{g}$)
to the terminal constraint $U(T)=U_{f}$, where the propagator $U$ solves the Cauchy problem $U(0)=I$ and
$\dot{U}=-\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left(U^{\dagger}H_{j}Ug\right)H_{j}U,\quad 0<t<1.$ (29)
Using Pontryagin’s maximum principle (see e.g. [5]), it can be shown that Problems ($\mathscr{P}_{u}$) and ($\mathscr{P}_{g}$) are equivalent [15] and admit solutions. The optimal time $T$ in the first problem corresponds in the second problem to the norm of the projection of an optimal $g$ on the space of controllable directions, that is $|Pg|$.
@@ C.2 Differential of the end-point mapping and its adjoint operator
In the framework of Problem ($\mathscr{P}_{g}$), the state $U:[0,1]\to SU(d)$ solves a non-linear equation with feedback control determined by $g\in\operatorname{i}\mathfrak{su}(d)$. More precisely we are interested in the Cauchy problem
$U(0)=I,\quad\dot{U}$ $=-\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(U^{\dagger}H_{j}Ug)H_{j}U,\quad t\in(0,1).$ (30)
Let $E:\operatorname{i}\mathfrak{su}(d)\to SU(d)$ denote the end-point mapping, defined by $E(g)=U(1)$ where $U$ is the solution of (30) with parameter $g$. The derivative $E^{\prime}$ of this mapping is needed to implement a first order optimization routine in the so-called optimize-then-discretize framework. We now focus on computing this derivative.
Equip the space of traceless Hermitian matrices $\operatorname{i}\mathfrak{su}(d)$ and $\mathbb{C}^{d\times d}$ with the inner-product $\langle a,b\rangle=\operatorname{Re}\operatorname{Tr}(ab^{\dagger})$. Then at any point $g$, the derivative $E^{\prime}(g):\operatorname{i}\mathfrak{su}(d)\to\mathbb{C}^{d\times d}$ and its adjoint $E^{\prime}(g)^{\dagger}:\mathbb{C}^{d\times d}\to\operatorname{i}\mathfrak{su}(d)$ are obtained as follows. As a shorthand notation, we write $h_{j}(t)=U(t)^{\dagger}H_{j}U(t)$.
Lemma C.4 (Jacobian-vector product [15]).
Given a perturbation $\delta g\in\operatorname{i}\mathfrak{su}(d)$, the resulting perturbation $\delta U(1)$ of the end-point $U(1)$ of the curve $U(t)$ is given by $E^{\prime}(g)\delta g=-\operatorname{i}U(1)w(1)$ where $w(t)\in\operatorname{i}\mathfrak{su}(d)$ satisfies
$w(0)=0,\quad\dot{w}=\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left(h_{j}\delta g-2\operatorname{i}gh_{j}w\right)h_{j},\quad t\in(0,1)$ (31)
Lemma C.5 (Vector-jacobian product [15]).
Given a linear form $S\in\mathbb{C}^{d\times d}$, the corresponding form $R=E^{\prime}(g)^{\dagger}S$ in $\operatorname{i}\mathfrak{su}(d)$ is given by
@@
$v(1)=\operatorname{i}U(1)^{\dagger}S,\quad\dot{v}=-2\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(h_{j}v^{\dagger})h_{j}g,\quad t\in(0,1)$ (33)
We refer to [15, Lemma 4.2, 4.3] for the details and a more general formula including a drift Hamiltonian and additional control variables. The calculation relies on the so-called adjoint-state method which we briefly outline and illustrate on a concrete example.
@@ C.2.1 The adjoint state method
Calculating the cost function $J(g)$ involves the implicit solution $U(t)$, to an equation of the form $F(U(\cdot),g)=0$ with parameter $g$. Loosely speaking implicit means that finding the solution $U$ given $g$ is a challenging computational task for which convenient closed form expressions cannot be exploited. In particular finding the gradient of $J$ is challenging because it is not given directly by a formula. The method is often a computationally efficient way to overcome this issue. At the core of the adjoint-state method lies the implicit function theorem and duality.
@@
Suppose that the vector parameter $g$ has $n$ real coordinates and that the cost function is
$J(g)=\ell(x(g))$
where $x(g)\in\mathbb{R}^{N}$ is defined as the unique solution to an implicit equation $F(x(g),g)=0$. Then in response to a perturbation $\delta g$ of the parameter $g$, the first variation of $J$ is
$\delta J=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}J(g+\varepsilon\delta g)\right|_{\varepsilon=0}=\langle\nabla\ell(x(g)),\delta x\rangle$ (34)
where $\delta x$ is the first variation of $x(g)$ given by
$\delta x=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}x(g+\varepsilon\delta g)\right|_{\varepsilon=0}$ (35)
We are evaluating a linear form on the implicit solution to a linear equation depending on $\delta g$. To relate $\delta x$ to $\delta g$ we match the first order variations in the equation $F(x(g),g)=0$. This implies that $\delta x$ solves a linear system of equation
$A\delta x+B\delta g=0$ (36)
where $A=\partial_{x}F(x(g),g)$ and $B=\partial_{g}F(x(g),g)$ are the jacobians of $F$ with respect to each arguments. The goal is to rewrite Equation (34) in the form $\delta J=\langle h,\delta g\rangle$ where the vector $h\in\mathbb{R}^{n}$ will be the gradient we seek. This is what enables the introduction of an adjoint state $v\in\mathbb{R}^{N}$ solution to the adjoint linear system
$A^{\dagger}v=\nabla\ell(x(g))$
@@
$\delta J=\langle x(g),\delta x\rangle=\langle A^{\dagger}v,\delta x\rangle=\langle v,A\delta x\rangle=\langle v,-B\delta g\rangle$
@@
$\boxed{\nabla J(g)=-B^{\dagger}v}$
which depends on $g$ via $B$ and the adjoint state $v$.
@@
$F(x,g)=\begin{pmatrix}\arctan(x_{1})-g_{1}\\ x_{2}-(x_{1}+g_{2})^{2}\end{pmatrix}$.
If $|g_{1}|<\pi/2$ then the vector $x(g)$ solution to $F(x,g)=0$ is given by
$x(g)=\begin{pmatrix}\tan g_{1}\\ (\tan g_{1}+g_{2})^{2}\end{pmatrix}$
This means that $J(g)=\tan(g_{1})^{2}/2+(\tan(g_{1})+g_{2})^{4}/2$ and so
$\partial_{g_{1}}J(g)$ $=(1+\tan(g_{1})^{2})\left(\tan(g_{1})+2(\tan(g_{1})+g_{2})^{3}\right),$ (37)
$\partial_{g_{2}}J(g)$ $=2(\tan(g_{1})+g_{2})^{3}.$ (38)
Let us obtain the same expression for $\nabla J(g)$ using the adjoint state method described above. With the shorthand notation $y=x(g)$, the coordinates of $y$ are $y_{1}=\tan g_{1}$ and $y_{2}=\tan g_{1}+g_{2}$ and so we have
$\nabla\ell(y)=y,\quad A=\begin{pmatrix}(1+y_{1}^{2})^{-1}&0\\ -2(y_{1}+g_{2})&1\end{pmatrix},\quad B=\begin{pmatrix}-1&0\\ 0&-2(y_{1}+g_{2})\end{pmatrix}$
Therefore the adjoint state $v$ has coordinates $v_{1}=(y_{1}+2y_{1}y_{2}+2g_{2}y_{2})(1+y_{1}^{2})$ and $v_{2}=y_{2}$. Accordingly the gradient $\nabla J(g)$ should be given by
$\nabla J(g)=\begin{pmatrix}v_{1}\\ 2(y_{1}+g_{2})v_{2}\end{pmatrix}$ (39)

### 10 Apr 2026 v2

@@ IV Conclusion
This work deals with the numerical implementation of a proof of concept to design smooth control pulses using a parametrization derived from the Pontryagin maximum principle. The corresponding method, coined MAGICARP, is applied to the implementation of quantum gates on spin qudit systems while negelecting the drift term. We demonstrated that, relative to the GRAPE method, the MAGICARP can yield significantly better execution times for spin systems up to dimension sixteen, and all the more as the dimension of the system is large. The method was showcased on two single molecule magnets, the $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$ molecules, taking into account some of their specific constraints.
@@
    Experimentally test the effectiveness of the pulses derived from the MAGICARP.
@@
    Extend the MAGICARP parametrization to closed system involving a drift Hamiltonian, and open systems.
@@
    Understand which optimizations and parametrizations could power up the MAGICARP, for instance by finding a clever gradient descents algorithm.
@@
    Relate the connectivity of the graph between the system’s energy levels to the efficiency of the method in terms of execution times.
@@ A.1 Importance of the graph
The graph on which simulations are taken are ones like the double decker from fig. 2a, linearly coupled graphs, and the one of the triple decker from fig. 2b. If we assume different graphs, possible in other physical molecules and platforms, we could have different time implementations for the same target gates and Hilbert space dimension. A most simple configuration is the linear graph where each level is only coupled to its nearest neighbours. It’s the congifuration of the fig. 1a. Another extremal configuration is the complete graph where each level is coupled to all the others. We can test how the GRAPE method and the MAGICARP take advantage of the growing connectivity of the graph. We will plot how the MAGICARP and the GRAPE method compare to themselves when runned over a complete graph rather than on a linearly coupled one in fig. 7. Note how in both case of fig. 7b the smallest execution time found is increasingly improved by the fact of using a complete graph instead of a linear one (thus having more available controls) with the size of the Hilbert space. In the case of fig. 7a the MAGICARP is slightly helped by the use of a complete graph instead of a linear one (gates are in median around 20% faster in all dimension) but is not as much of an improvement as for GRAPE. Thus sparsely connected graph seem already convenient for the MAGICARP.
Figure 7: Figure comparing the median or the minimum time to make at least 100 gates with a precision of 1e-4 for the MAGICARP and the GRAPE method on complete graphs with 3 to 9 nodes with themselves on linear graphs. For each optimization method, the quantity shown is the ratio between the runtime on a linear graph and the runtime on a complete graph, effectively describing how each method is faster when being run on a complete graph instead of a linear one. (a) Ratio of the median execution times for 100 random Haar gates. (b) Ratio of the minimum execution times for a QFT gate.
@@ A.2 Importance of a $\sigma_{z}$ control
Figure 8: Figure comparing the median or the minimum time to make at least 100 gates with a precision of 1e-4 for the MAGICARP and the GRAPE method on complete and linear graphs with 3 to 9 nodes with themselves when allowing for a control on $\sigma_{z}$. For each optimization method, the quantity shown is the ratio between the runtime with only $\sigma_{x}$ and $\sigma_{y}$ controls and the runtime with all three controls on each edge of the graph the method has been run on. It effectively describes how each situation (method + graph it is run on) is faster when being run while allowing a control on $\sigma_{z}$ instead of just $\sigma_{x}$ and $\sigma_{y}$. The red curves correspond to the aforementioned ratio for the GRAPE method run on a linear graph, the orange ones for the GRAPE method run on a complete graph, the green ones for the MAGICARP run on a lienar graph, and the cyan ones for the MAGICARP run on a complete graph. The x-axis describes the Hilbert space dimension the optimization methods have been run on. (a) Ratio of the median execution times for 100 random Haar gates. (b) Ratio of the minimum execution times for a QFT gate.
One can note that adding a control on $\sigma_{z}$ is not improving much the median solutions found on random gates in fig. 8a or the minimum solutions found in fig. 8b for the specific QFT gate in all situations but in the case of the MAGICARP method run on a complete graph, the cyan curve of fig. 8a, especially when the dimension grows. It applies only for the median execution time since the same cyan curve in fig. 8a is not affected. It would be interesting to understand further why adding a $\sigma_{z}$ control gives a particular advantage to the MAGICARP on a strongly connected graph.
@@ B.1 Improvement of the GRAPE with the target fidelity
The GRAPE method gives better execution times according to the target fidelity it has too reach.
@@
It is understandable since being able to be further away from the target obviously relieves a constraint on the pulses. We chose to compare execution times at 1e-4 of target fidelity for the GRAPE when comparing to other algorithms for multiple reasons. First, this is already a fidelity allowing for making a short algorithm without being bothered by this lack of precision of the gates while in many platforms noise is already inducing an error of this order [28]. Second, Maintaining the fidelities of pulses found by the MAGICARP to be strictly lower than 1e-4 proved difficult with the gradient we used. Third, at lower fidelities than 1e-4 GRAPE is not giving significantly slower pulses than for this value of the target fidelity as shown infig. 9.
@@ B.2 Improvement of the GRAPE with the number of controls
The GRAPE method gives better execution times according to the number of piece-wise amplitude steps it can use, though up to a limit. In all this article, we chose to compare execution times with $10(\text{dim}+1)$ piece-wise amplitude steps for each controls, with dim being the dimension of the Hilbert space at hand. It is a large enough number of amplitude to not see improvements by taking a larger amount of piece-wise amplitude steps for each controls as shown infig. 10.
@@ Appendix C Numerical framework
Our objectives in this section is twofold. We provide closed-form expressions for the derivative of the map that associates to every control parameter $g$ the final propagator of Schrödinger’s equation. We outline its derivation which relies on the so-called adjoint state method. This provide the tools needed to implement any first-order optimization routine and possibly to adapt the $M$ parametrization of controls to another problem. On the other hand, we propose our implementation of the natural gradient descent, an optimization routine used to produce some of the figures in this paper. Julia and Python codes are available on the Github repositories Magicarp and pyMagicarp.
@@ C.1 Two equivalent optimal control problems in the driftless case
The third Assumption C.3 is common in the practical coherent control of quantum systems because the control Hamiltonians often do not linearly span the set of tangent vectors to the special unitary group at the identity matrix. The Lie algebra rank condition is about iterated commutators of the control Hamiltonians. When it holds, combining the available control directions provides a family of tangent vectors that is sufficiently rich to move in any admissible direction. Intuitively this is because to second order in $t$ close to zero, $e^{-tA}e^{-tB}e^{tA}e^{tB}=I+t^{2}[A,B]+o(t^{2})$. All in all Assumption C.3 is meant to guarantee that any desired special unitary gate is reachable by means of an appropriate sequence of constant pulses.
@@
Let $P$ stand for the orthogonal projector onto the real linear span of $H_{1},\ldots,H_{m}$ for the Frobenius (Hilbert-Schmidt) norm $M\mapsto\operatorname{Tr}(MM^{\dagger})^{1/2}$. Consider the following two optimal control problems. On the one hand
@@
subject to the pointwise constraint $\sum_{j=1}^{m}u_{j}(t)^{2}\leq 1$ and to the terminal constraint $U(T)=U_{f}$, where the propagator $U$ solves the Cauchy problem $U(0)=I$ and $\dot{U}(t)=-\operatorname{i}\sum_{j=1}^{m}u_{j}(t)H_{j}U(t)$ for $0<t<T$.
@@
$\inf_{M\in\operatorname{i}\mathfrak{su}(d)}|P(M)|$ ($\mathscr{P}_{M}$)
to the terminal constraint $U(1)=U_{f}$, where the propagator $U$ solves the Cauchy problem $U(0)=I$ and
$\dot{U}(s)=-\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left(U(s)^{\dagger}H_{j}U(s)M\right)H_{j}U(s),\quad 0<s<1.$ (44)
Using Assumption C.3 and rescaling, we can build an admissible control for Problem ($\mathscr{P}_{u}$) which steers the identity to $U_{f}$ while satisfying the experimental constraint $\sum_{j=1}^{m}u_{j}(t)^{2}\leq 1$. This relies on the Chow-Rashevskii theorem or more specifically [2, Corollary 3.1]. Moreover it can be shown that an optimal solution exists and that, using the maximization condition from Pontryagin’s maximum principle (see e.g. [8]), any normal extremal is admissible for Problem ($\mathscr{P}_{M}$) with
$T=|P(M)|,\quad\text{ and }\quad u_{j}(t)=\operatorname{Re}\operatorname{Tr}\left(U(s)^{\dagger}H_{j}U(s)\frac{M}{|P(M)|}\right),\quad\text{where}\quad\quad Ts=t.$ (45)
The time $T$ in the first problem corresponds in the second problem to the norm of the projection of $M$ onto the space of controllable directions, that is $|P(M)|$.
Remark 2 (Restriction to normal extremals).
Not all Pontryagin extremal are normal and hence, it is not true in general that every optimal control $u$ of Problem ($\mathscr{P}_{u}$) is of the form given by Equation (45). Depending on the choice of control Hamiltonians, one can prove that any abnormal extremal, which is not normal at the same time, cannot be optimal and hence can be disregarded. This occurs for instance in the $K+P$ setting. The proof given in [7, Appendix C] relies on the Cartan decomposition of Lie groups and the Goh necessary conditions derived in [1].
@@ C.2 Differential of the end-point mapping and its adjoint operator
In the framework of Problem ($\mathscr{P}_{M}$), the state $U:[0,1]\to SU(d)$ solves a non-linear equation with feedback control determined by $M\in\operatorname{i}\mathfrak{su}(d)$. More precisely we are interested in the Cauchy problem
$U(0)=I,\quad\dot{U}$ $=-\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(U^{\dagger}H_{j}UM)H_{j}U,\quad t\in(0,1).$ (46)
Let $E:\operatorname{i}\mathfrak{su}(d)\to SU(d)$ denote the end-point mapping, defined by $E(M)=U(1)$ where $U$ is the solution of (46) with parameter $M$. The derivative $E^{\prime}$ of this mapping is needed to implement a first order optimization routine in the so-called optimize-then-discretize framework. We now focus on computing this derivative.
Equip the space of traceless Hermitian matrices $\operatorname{i}\mathfrak{su}(d)$ and $\mathbb{C}^{d\times d}$ with the inner-product $\langle a,b\rangle=\operatorname{Re}\operatorname{Tr}(ab^{\dagger})$. Then at any point $M$, the derivative $E^{\prime}(M):\operatorname{i}\mathfrak{su}(d)\to\mathbb{C}^{d\times d}$ and its adjoint $E^{\prime}(M)^{\dagger}:\mathbb{C}^{d\times d}\to\operatorname{i}\mathfrak{su}(d)$ are obtained as follows. As a shorthand notation, we write $h_{j}(t)=U(t)^{\dagger}H_{j}U(t)$.
Lemma C.4 (Jacobian-vector product).
Given a perturbation $\delta M\in\operatorname{i}\mathfrak{su}(d)$, the resulting perturbation $\delta U(1)$ of the end-point $U(1)$ of the curve $U(t)$ is given by $E^{\prime}(M)\delta M=-\operatorname{i}U(1)w(1)$ where $w(t)\in\operatorname{i}\mathfrak{su}(d)$ satisfies
$w(0)=0,\quad\dot{w}=\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left(h_{j}\delta M-2\operatorname{i}Mh_{j}w\right)h_{j},\quad 0<t<1$ (47)
Lemma C.5 (Vector-jacobian product).
Given a linear form $S\in\mathbb{C}^{d\times d}$, the corresponding form $R=E^{\prime}(M)^{\dagger}S$ in $\operatorname{i}\mathfrak{su}(d)$ is given by
@@
$v(1)=\operatorname{i}U(1)^{\dagger}S,\quad\dot{v}=-2\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(h_{j}v^{\dagger})h_{j}M,\quad 0<t<1$ (49)
We postpone the proofs of these results to Section C.2.2. The calculation relies on the so-called adjoint-state method which we briefly outline and illustrate on a concrete example.
@@ C.2.1 The adjoint state method
Calculating the cost function $J(M)$ involves the implicit solution $U(t)$, to an equation of the form $F(U(\cdot),M)=0$ with parameter $M$. Loosely speaking implicit means that finding the solution $U$ given $M$ is a challenging computational task for which convenient closed form expressions cannot be exploited. In particular finding the gradient of $J$ is challenging because it is not given directly by a formula. The method is often a computationally efficient way to overcome this issue. At the core of the adjoint-state method lies the implicit function theorem and duality.
@@
Suppose that the vector parameter $M$ has $n$ real coordinates and that the cost function is
$J(M)=\ell(x(M))$
where $x(M)\in\mathbb{R}^{N}$ is defined as the unique solution to an implicit equation $F(x(M),g)=0$. Then in response to a perturbation $\delta M$ of the parameter $M$, the first variation of $J$ is
$\delta J=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}J(M+\varepsilon\delta M)\right|_{\varepsilon=0}=\langle\nabla\ell(x(M)),\delta x\rangle$ (50)
where $\delta x$ is the first variation of $x(M)$ given by
$\delta x=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}x(M+\varepsilon\delta M)\right|_{\varepsilon=0}$ (51)
We are evaluating a linear form on the implicit solution to a linear equation depending on $\delta M$. To relate $\delta x$ to $\delta M$ we match the first order variations in the equation $F(x(M),M)=0$. This implies that $\delta x$ solves a linear system of equation
$A\delta x+B\delta M=0$ (52)
where $A=\partial_{x}F(x(M),M)$ and $B=\partial_{M}F(x(M),M)$ are the jacobians of $F$ with respect to each arguments. The goal is to rewrite Equation (50) in the form $\delta J=\langle h,\delta M\rangle$ where the vector $h\in\mathbb{R}^{n}$ will be the gradient we seek. This is what enables the introduction of an adjoint state $v\in\mathbb{R}^{N}$ solution to the adjoint linear system
$A^{\dagger}v=\nabla\ell(x(M))$
@@
$\delta J=\langle x(M),\delta x\rangle=\langle A^{\dagger}v,\delta x\rangle=\langle v,A\delta x\rangle=\langle v,-B\delta M\rangle$
@@
$\boxed{\nabla J(M)=-B^{\dagger}v}$
which depends on $M$ via $B$ and the adjoint state $v$.
@@
$F(x,M)=\begin{pmatrix}\arctan(x_{1})-M_{1}\\ x_{2}-(x_{1}+M_{2})^{2}\end{pmatrix}$.
If $|M_{1}|<\pi/2$ then the vector $x(M)$ solution to $F(x,M)=0$ is given by
$x(M)=\begin{pmatrix}\tan M_{1}\\ (\tan M_{1}+M_{2})^{2}\end{pmatrix}$
This means that $J(M)=\tan(M_{1})^{2}/2+(\tan(M_{1})+M_{2})^{4}/2$ and so
$\partial_{g_{1}}J(M)$ $=(1+\tan(M_{1})^{2})\left(\tan(M_{1})+2(\tan(M_{1})+M_{2})^{3}\right),$ (53)
$\partial_{g_{2}}J(M)$ $=2(\tan(M_{1})+M_{2})^{3}.$ (54)
Let us obtain the same expression for $\nabla J(M)$ using the adjoint state method described above. With the shorthand notation $y=x(M)$, the coordinates of $y$ are $y_{1}=\tan M_{1}$ and $y_{2}=\tan M_{1}+M_{2}$ and so we have
$\nabla\ell(y)=y,\quad A=\begin{pmatrix}(1+y_{1}^{2})^{-1}&0\\ -2(y_{1}+M_{2})&1\end{pmatrix},\quad B=\begin{pmatrix}-1&0\\ 0&-2(y_{1}+M_{2})\end{pmatrix}$
Therefore the adjoint state $v$ has coordinates $v_{1}=(y_{1}+2y_{1}y_{2}+2M_{2}y_{2})(1+y_{1}^{2})$ and $v_{2}=y_{2}$. Accordingly the gradient $\nabla J(M)$ should be given by
$\nabla J(M)=\begin{pmatrix}v_{1}\\ 2(y_{1}+M_{2})v_{2}\end{pmatrix}$ (55)